Two kinds of bacteria are found in a sample of tainted food. It is found that the population size of type 1, N1 and of type 2, N2 satisfy the equation

dN/dt=-0.2/N1

dN/dt=0.6/N2

N1 is equal to 2000 at time equal to zero, while N2 is equal to 20 at time equal to zero.

Then the population sizes are equal (N1 = N2) at what time? (4 decimal places)


1
Expert's answer
2020-12-22T19:54:19-0500

Let us solve the equation "\frac{dN_1}{dt}=-\frac{0.2}{N_1},\ \ N_1(0)=2000."


"N_1dN_1=-0.2dt"


"\int N_1dN_1=-\int 0.2dt"


"\frac{(N_1)^2}{2}=-0.2t+C_1"


If "N_1(0)=2000", then "\frac{2000^2}{2}=C_1", and thus "C_1=2,000,000".


So, "(N_1)^2=-0.4t+4,000,000."



Let us solve the equation "\frac{dN_2}{dt}=\frac{0.6}{N_2},\ \ N_2(0)=20."


"N_2dN_1=0.6dt"


"\int N_2dN_2=\int 0.6dt"


"\frac{(N_2)^2}{2}=0.6t+C_2"


If "N_2(0)=20", then "\frac{20^2}{2}=C_2", and thus "C_2=200".


So, "(N_2)^2=1.2t+400"


If "N_1=N_2", then "(N_1)^2=(N_2)^2" , and thus "-0.4t+4,000,000=1.2t+400". It follows that "1.6t=3,999,600", and therefore, "t=2,499,750."


Answer: "t=2,499,750."